OpenStudy (resal):

If f(x) = 2x2 + 3x and g(x) = x – 2, what is (f + g)(2)?

OpenStudy (tkhunny):

That notation means simply to add the two functions together. (f+g)(2) = f(2) + g(2) Evaluate each and add.

OpenStudy (kittiwitti1):

Elaborating on tkhunny explanation...\[\color{red}{f(x)=2x^2+3x},~\color{blue}{g(x)=x-2}\]therefore...\[(\color{red}{f}+\color{blue}{g})(x)=\color{Red}{f(x)}+\color{Blue}{g(x)=\color{Red}{2x^2+3x}+\color{Blue}{x-2}}\]now simplify...\[2x^2+4x-2\]

OpenStudy (kittiwitti1):

So you have \((f+g)(x)=2x^2+4x-2\), now you want to find \((f+g)(2)\) so input \(x=2\)

OpenStudy (jerry45):

ok.....